The following information is taken from the accounts of Latta Company. The entries in the T-accounts are summaries of the traRequired 1 Required 2 Required 3 Identify reasons for entries (a) through (d). Reason Entry (a) (b) (c) (d)Required 1 Required 2 Required 3 Assume that the underapplied or overapplied overhead is closed to cost of Goods Sold. PreparRequired 1 Required 2 Required 3 Assume that the underapplied or overapplied overhead is closed proportionally to Work in Pro
Answer #1

1. Reasons for Entries:

a) Manufacturing Overhead which has been paid during the year has been 488448.

b) Out of Manufacturing Overhead of 488448 , only amount of 407040 has belong the current year, that is why that amount has been tranfered to Work in progress account and balance of 81408 belongs to next year and will be tranfer to WIP Account Accordingly.

c) Tranfer of 758000 from WIP Account to Finished Goods Account : The closing balance of WIP Account is 52000 , out of which 48% has been Manufacturing Overhead as per the Overhead absorption rate(given in question) ie 24960 from manufacturing overhead tranfered to WIP Acount(407040) has been included in closing balance and balance of Manufacturing overhead of 382080(407040-24960) along with opening balance of WIP Account 9960 and other expenses of 365960(302000+91000-27040) which comes out to be 758000(382080+9960+365960) has been tranfered to Finished Goods account. However Overhead has been under absorbed.

b) Tranfer of 664000 from Finished Goods Account to cost of goods sold account : The closing balance of Finished Goods Account is 132000 , out of which 48% has been Manufacturing Overhead as per the Overhead absorption rate(given in question) ie 63360 from manufacturing overhead tranfered to finished Goods Acount(382040) has been included in closing balance and balance of Manufacturing overhead of 318720(382040-63360) along with opening balance of Finished goods Account 38000 and other expenses of 307280(3659650+9960 from WIP Account- 68640 from closing of FG) which comes out to be 664000(318720+38000+307280) has been tranfered to Finished Goods account.

2. Journal Entry when Overapplied or underapplied Overhead has been closed to Cost of Goods sold:

Manufacturing Overhead Recovery Rate = Overhead / Other Expenses

(Based on Closing balance of WIP)

= 24960(given)/ 27040(52000-24960)

=92.3077%

Recoverable OH on WIP account = 402940*92.3077% = 371963

Less : REcovered Applied on WIP Account = 402960

Overapplied Overhead on WIP Account = 30996.92

Recoverable OH on FG account = 372320*92.3077% = 343680

Less : REcovered Applied on FG Account = 291680

Underapplied Overhead on FG Account = 28640

General Journal
Sr. No Particulars Debit Credit
1 Cost of Goods Sold A/c Dr 30996.92
To Work in Progress Account 30996.92
(Being Overapplied and underapplied overhead closed to Cost of Goods sold account)
2 Finished Goods A/c Dr 28640
To Cost of Goods Sold A/c 28640
(Being Overapplied and underapplied overhead closed to Cost of Goods sold account)

3. Journal Entry when Overapplied or underapplied Overhead has been closed proportionally to WIP, FG, COGS:

General Journal
Sr. No Particulars Debit Credit
1 Cost of Goods Sold A/c Dr 2356.92
Finished Goods A/c Dr 28640
To Work in Progress Account 30996.92
(Being Overapplied and underapplied overhead closed to Cost of Goods sold account)